outline:

introduction
stokes theorem
relevance of free-surface
importance of integral
completeness
individuality of sources
time zero: Papanicolau freq. domain smoothing=time window
current goals
source location
statics correction
imaging
datum-correlate
complicated initial conditions
imprecision in rigorous theory effects on direct migration
multiple sources
insufficient sources
undersampled wavefields
CASC
introduction
location
previous work
data
bandwidth
preprocessing
imaging
image modes
imaging vs. parameter inversion
v(x,z) migration
CIG's and velocity information
direct migration
discussion/implications

Valhall
introduction
location
hardware description
timing/extent
25 hours Feb. 15 2004
12 hours Jan. 2005
time processing
correlation vs. deconvolution
stability
number of time windows to reveal signal
identification of signal
gathers
         off-array location
         identification
                 model possibilities
                           point source
                           reflection from a plane
                           2 vs. 3 vs. 5/7 raypaths
                 location
                           inversion
                           precision
                           time stability
cmp sorted data
         poorness
         goodness
direct migration
active migration benchmark
parameter selection
time window processing
         buildup
         changes over time
         depth stacked results
time-stacked processing
correlations w/ active migration

Moss Beach experiment
introduction
         location
         hardware
         data
         active migration benchmark
time window stability
deconvolution
anomaly identification
analysis/identification of areas for improvement
Discussion/implications
incomplete integrals
         incorrecte kinematics
         illumination holes
interpretation data volume
         time processing
         depth processing
potential goodness
         time v. depth
addressing badness
         nature imposed
         equipment/technique remediable
Multiple prediction as an imaging experiment
commutability of extrapolation and convolution
SRMP v. IS-SRMP
Sigsbee
Discussion/implications